# Invar-RAG: Invariant LLM-aligned Retrieval for Better Generation

### v1

@@ Abstract
Retrieval-augmented generation (RAG) has shown its impressive capability of providing reliable answer predictions and addressing severe hallucination problems. A typical RAG implementation adopts powerful retrieval models to extract external information and leverage large language models (LLMs) to generate corresponding answers. Different with that, recent LLM-based retrieval has raised much attention because it brings substantial improvements in information retrieval (IR) via LLMs’ vigorous semantic understanding capability. However, directly applying LLM to RAG systems remains certain challenges. This may cause feature locality problems since massive parametric knowledge impedes the effective usage of the global information among all corpus, e.g., a LLM-based retriever usually inputs the summary of documents instead of the whole documents. Moreover, various tasks pre-trained in LLMs induce severe variance, which further weakens its performance as the retriever. To address these issues, we propose a novel two-stage fine-tuning architecture called Invar-RAG. In the retrieval stage, a LLM-based retriever is constructed by integrating a LoRA-based representation learning to address the feature locality problem. To justify and consolidate this retrieval’s performance, two patterns (i.e., invariant and variant patterns) and an invariance loss are also developed to alleviate the variance in LLM. Moreover, in the generation stage, a meticulously designed fine-tuning method is devised to improve our LLM for accurate answer generation based on the retrieved information. Experimental results demonstrate that Invar-RAG significantly outperforms existing baselines across three Open-domain Question Answering (ODQA) datasets. The code is available in Supplementary Material to ease reproducibility.
1 Introduction
Over the past decade, large language models (LLMs) have demonstrated promising capability in processing natural language Bajaj et al. (2016). Owing to the vast amount of knowledge encoded in their internal parameters, LLMs such as GPT Izacard & Grave (2020) and LLaMa Yao et al. (2014) have demonstrated remarkable performance on various downstream tasks, including Open-domain Question Answering (ODQA) Zhu et al. (2021), Reading Comprehension Cheng et al. (2023). However, the fixed parametric knowledge of LLMs has hindered the further applications of LLMs and made them prone to errors (hallucination Wongsuphasawat et al. (2012) and factual errors Fan et al. (2024)).
@@ 1 Introduction
To overcome the limitations of parametric knowledge, one promising approach is Retrieval-Augmented Generation (RAG) Fan et al. (2024); Lewis et al. (2020). Compared to relying solely on parametric knowledge, RAG enables LLMs to use retrievers to access relevant information from external knowledge sources, enhancing their question-answering abilities. Among the two components of RAG, current methods primarily focus on optimizing the retriever to return more relevant documents due to the high cost of fine-tuning and black-box LLM APIs. Previous retrievers leveraged deep learning technology (e.g., dense retrieval Zhao et al. (2024)) to encode the text representations from the lexical space into the high-dimensional latent space, allowing them to model more complex semantic relationships between queries and corpora. However, the separation between the retriever and generation model has hindered their full integration, limiting their compatibility in downstream applications . Some advanced RAG systems, such as RA-DIT Lin et al. (2023b), have adopted joint training mechanisms that fine-tune both the retriever and the generation model for better alignment. However, this approach is impractical due to the need for frequent fine-tuning and fails to utilize the LLMs’ semantic understanding capabilities during the retrieval stage.
Consequently, generative retrieval (GR), also known as LLM-based retrieval, leverages the parametric memory of generative models to directly generate document identifiers (DocIDs) Li et al. (2024), which has aroused much attention. By memorizing the documents as the parametric knowledge of LLM, this kind of method breaks the limitations of traditional IR in terms of document granularity and simple relevance matching Nguyen & Yates (2023), offering more flexibility and creativity, thus better meeting practical needs. However, two severe problems hinder the current LLM-based retrieval. 1)Feature Locality: LLM-based retrieval normally adopt language models to learn the mapping from queries to the relevant document DocIDs. However, these DocIDs actually can not fully represent the global information of the passages. Meanwhile, directly feeding the whole passages into LLM is costly and infeasible, causing a trade-off between effectiveness and efficiency. 2)Retrieval Variance: Due to the inherent generative inconsistency property of large language models , current LLM-based retrieval may generate unforeseeable variances, especially when the input query or the size of context varies, directly results in undesirable and vulnerable performance which may not be preferred.
@@
Considering the problems mentioned above and better leveraging the capability of LLMs, we propose a fully LLM-based architecture with a two-stage fine-tuning method called Invar-RAG, as illustrated in Figure 1. In the retrieval stage, our approach initializes the pre-trained LLaMA Yao et al. (2014) as the backbone and follows the bi-encoder architecture in DPR Karpukhin et al. (2020) to construct our retriever. Compared to normal GR methods which need an iterative process of encoding and decoding, we introduce a component called LLM-aligned Retrieval. It first represents the input query and corpora into high-dimension space using a small language model (MiniLM) Wang et al. (2020), then introduces a new loss function constructed by KL-divergence to align the coarse query-documents pairs representation to the LLM’s representation space. This allows the retriever to leverage the rich prior knowledge of LLM, typically addressing the feature locality caused by only feeding DocIDs to LLM. Moreover, based on the initial objective of our LLM-aligned Retrieval, we introduce the invariance loss to overcome the variance in the retrieval stage. By recognizing the invariant pattern that contributes the most to the performance and gradually forcing the model to rely on the invariant pattern, we can avoid the unforeseeable variances in practice and enhance the robustness of our RAG system. Finally, in the generation stage, we freeze the weights we fine-tuned before and optimize the generation function to allow the LLM to give correct answers to the retrieved documents. Our contributions are summarized as follows:
@@
Figure 1: Overview of proposed Invar-RAG.
2 Methodology
@@
2.1 Overall Framework of Invar-RAG
@@
2.2 Retrieval Stage
Architecture. For the Retriever architecture, we follow the approach of previous work Ma et al. (2024), using the bi-encoder architecture from DPR Karpukhin et al. (2020), but replacing the backbone model with LLaMA Yao et al. (2014). Considering the efficiency, we first compute the vector embedding of a document ${d}^{R}_{i}\in\mathcal{D}^{R}$ as:
${V}_{r}({{d}^{R}_{i}})=\operatorname{Decoder}\left(`t_{1}␣t_{2}␣\cdots␣t_{k}^{\prime}\right)\left[-1\right]$ (1)
@@ 2.2 Retrieval Stage
$s_{raw}\left(V_{q},V_{r}({{d}_{i}^{R}})\right)$ $=V_{q}\cdot V_{r}({{d}_{i}^{R}})$ (2)
@@
LLM-aligned Retrieval. Building on the above structure, we enhance our LLM-based retriever’s ability to return more relevant documents. We introduce LLM-aligned retrieval with invariance loss in the retrieval stage, effectively addressing the aforementioned issues. Current alignment methods, such as RA-DIT Lin et al. (2023b), primarily focus on aligning the scoring functions between the retriever and the generator. However, the initial structural differences between novel retrievers (e.g., DRAGON+ Lin et al. (2023a)) and LLMs still impede further optimization of the overall RAG system. Therefore, we design a novel LLM-based retriever to resolve this issue. Unlike previous LLM-based GR methods Li et al. (2024), we no longer need to use DocIDs to retrieve relevant documents, which may cause the feature locality issue mentioned earlier. Instead, we adopt a fine-tuned alignment process that enables the LLM to perform representation learning. We leverage LoRA architecture Hu et al. (2021) to add additional adapter parameter $\theta_{\mathcal{R}}$ to our raw representation $V_{r}$, denoted as $V_{r}({{d}_{i}^{R}},\theta_{\mathcal{R}})$. The corresponding relevance score can then be re-normalized among top-k relevant chunks $\mathcal{D}^{{R}^{\prime}}\subset\mathcal{D}^{{R}}$ as:
$S^{r}_{\mathcal{R}}(V_{r}({{d}_{i}^{R}},\theta_{\mathcal{R}})|V_{q})=\frac{\exp{s_{raw}\left(V_{q},V_{r}({{d}_{i}^{R}},\theta_{\mathcal{R}})\right)}}{\sum_{{d_{i}^{{R}}}^{\prime}\in\mathcal{D}^{{R}^{\prime}}}\exp{s_{raw}\left(V_{q},V_{r}({{{d}_{i}^{R}}^{\prime}},\theta_{\mathcal{R}})\right)}}$ (3)
@@
For each document in the corpus, we need to compute $S^{r}_{\mathcal{R}}$ for $n$ times ($n$ represents the number of documents in $\mathcal{D}^{R}$) to rank the relevance scores. Consequently, the initial loss function for representation learning can then be defined by minimizing the KL-divergence Kim et al. (2021) of two relevance scores leveraging Eq. 3.
$\mathcal{L}_{rl}\left(\mathcal{D}^{R}\right)=\mathbb{E}_{d_{i}^{R}\in\mathcal{D}^{R}}\operatorname{KL}(S^{r}_{\mathcal{R}}(V_{r}({{d}_{i}^{R}},\theta_{\mathcal{R}})|V_{q})\|S^{r}_{\mathcal{R}}(V_{p}({{d}_{i}^{R}})|V_{q}))$ (4)
@@
Following the previous works Lin et al. (2023b); Ma et al. (2020), fine-tuning both encoders hurt the performance Bao et al. (2024), we only update a part of our initialized retriever, which is in charge of computing the query representation.
@@
Invariance Loss. To further enhance retrieval accuracy while maintaining robustness, we introduce invariance loss, building on our initial KL-divergence loss. Current refinement methods primarily rely on query rewriting He et al. (2016); Chan et al. (2024) or LLM generation Fan et al. (2024) to expand the search space and re-rank document chunks. However, they fail to recognize the effectiveness of different rewriting and generation procedures, directly resulting in the invariance problem. Specifically, we begin by rewriting the query and adjusting the context window to broaden the search space. Since not all refinement methods are effective, we identify invariant patterns to preserve retrieval performance while gradually incorporating weighted variant patterns to broaden the search space. To achieve this, we use the LSR score from LM-Supervised Retrieval Shi et al. (2023) to determine whether a document effectively enhances the LLM’s answer prediction capability. For a training sample $(q,y)$, where q and y respectively represent the input query and output result, we first define the output probability of LM as:
$p_{LM}(y|{V}(d_{i}^{R}\circ x))=\sum_{{d}^{R}_{i}\in\mathcal{D}^{R}}p_{LM}(y|{V}({d}^{R}_{i}\circ x))\cdot P_{\mathcal{R}}({d}^{R}_{i}|x)$ (5)
@@
$P_{LSR}(d_{i}|q,y)=\frac{\exp(p_{LM}(y|d_{i}\circ q)/\tau)}{\sum_{d_{i}^{\prime}\in{\mathcal{D}^{R}}^{\prime}}\exp(p_{LM}(y|d_{i}^{\prime}\circ q)/\tau)}\approx\frac{\exp(p_{LM}(y|d_{i}\circ q)/\tau)}{\sum_{d_{i}^{\prime}\in\mathcal{D}^{R}}\exp(p_{LM}(y|d_{i}^{\prime}\circ q)/\tau)}$ (6)
@@
$\mathcal{L}_{invar}(\mathcal{D}_{in})=\operatorname{Var}_{\mathcal{D}\subseteq\mathcal{D}_{var}}(\mathbb{E}_{d_{in}^{R}\in(\mathcal{D}_{in}\cup\mathcal{D})}\operatorname{KL}(P^{r}_{\mathcal{R}}(V_{r}(d_{in}^{R},\theta_{\mathcal{R}})|V_{q})\|P^{p}_{\mathcal{R}}(V_{p}(d_{in}^{R})|V_{q})))$ (7)
@@
$\min_{\theta_{R}}\mathcal{L}_{rl}+\lambda\mathcal{L}_{invar}$ (8)
@@
2.3 Generation Stage
In the generation stage, We followed the same architecture as in the retrieval stage for answer prediction. To improve the generative capability of LLM for leveraging the retrieved information better, followed by prior works Lin et al. (2023b); Shi et al. (2023), we adopt another LoRA adapter to fine-tune our model on different tasks. Specifically, for the same training sample $(x,y)$, we retrieve the top-$\overline{k}$ relevant document chunks $\mathcal{D}_{G}^{\prime}\subset\mathcal{D}_{G}$ by performing our model on retrieval task. For each retrieved chunk $d_{i}\in\mathcal{D}_{G}^{\prime}$, we design a special fine-tuning example by prepending it to the prompt as background information and create k independent instances for one original example: $\left\{(d_{i}\circ x,y)|i=1,\cdots,\overline{k}\right\}$. Then, following the previous work Lin et al. (2023b); Qi et al. (2020), we fine-tune the language model using the next-token prediction objective and minimize the loss as follows:
$\mathcal{L}(\mathcal{D}_{G}^{\prime})=-\sum_{i}\log p_{LM}(y|d_{i}\circ x)$ (9)

### v2

@@ Abstract.
Retrieval-augmented generation (RAG) has stood out as a highly promising technique in numerous domains due to its impressive capability of giving reliable answer predictions and addressing the severe hallucination problem. Current RAG systems employ powerful retrieval models to extract external information and leverage large language models (LLMs) as the generation model. Nevertheless, the differences in the architecture between the retriever and generation model have often resulted in notable ineffectiveness, posing a significant obstacle to retrieving and predicting the answers. Recently, LLM-based retrieval has demonstrated its exceptional effectiveness in information retrieval (IR), delivering substantial improvements via LLMs’ powerful semantic understanding capability. However, directly applying LLM to RAG systems poses certain challenges. Its massive parametric knowledge impedes the ability to leverage the global information of all the corpus, leading to a scene that current LLM-based retriever usually inputs the summary of documents instead of the whole documents, which causes the feature locality problem. Moreover, various tasks pre-trained in LLMs induce severe variance, which further weakens its performance as the retriever. To address these issues, we propose a novel two-stage fine-tuning architecture called Invar-RAG. In the retrieval stage, by introducing representation learning leveraging the LoRA structure, we construct a LLM-based retriever to address the feature locality problem. Moreover, to consolidate our designed retrieval’s performance, we define two patterns (invariant pattern and variant pattern) and propose the invariance loss to eliminate the variance in LLM to some extent. Then, for the generation stage, we devise another fine-tuning method to allow our LLM to better leverage the retrieved information to generate an accurate answer. Experimental results demonstrate that Invar-RAG significantly outperforms existing baselines across three Open-domain Question Answering (ODQA) datasets. Our implementation code is available in Supplementary Material to ease reproducibility.
LLM, Information Retrieval
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06
1. Introduction
@@
Over the past decade, large language models (LLMs) have demonstrated promising capability in processing natural language (Bajaj et al., 2016). Owing to the vast amount of knowledge encoded in their internal parameters, LLMs such as GPT (Izacard & Grave, 2020) and LLaMa (Yao et al., 2014) have demonstrated remarkable performance on various downstream tasks, including Open-domain Question Answering (ODQA) (Zhu et al., 2021), Reading Comprehension (Cheng et al., 2023). However, the fixed parametric knowledge of LLMs has hindered the further applications of LLMs and made them prone to errors (hallucination (Wongsuphasawat et al., 2012) and factual errors (Fan et al., 2024)). To overcome the limitations of parametric knowledge, one promising approach is Retrieval-Augmented Generation (RAG) (Fan et al., 2024; Lewis et al., 2020). Compared to relying solely on parametric knowledge, RAG enables LLMs to use retrievers to access relevant information from external knowledge sources, enhancing their question-answering abilities. Among the two components of RAG, current methods primarily focus on optimizing the retriever to return more relevant documents due to the high cost of fine-tuning and black-box LLM APIs. Previous retrievers leveraged deep learning technology (e.g., dense retrieval (Zhao et al., 2024)) to encode the text representations from the lexical space into the high-dimensional latent space, allowing them to model more complex semantic relationships between queries and corpora. However, the separation between the retriever and generation model has hindered their full integration, limiting their compatibility in downstream applications . Some advanced RAG systems, such as RA-DIT (Lin et al., 2023b), have adopted joint training mechanisms that fine-tune both the retriever and the generation model for better alignment. However, this approach is impractical due to the need for frequent fine-tuning and fails to utilize the LLMs’ semantic understanding capabilities during the retrieval stage. Consequently, generative retrieval (GR), also known as LLM-based retrieval, leverages the parametric memory of generative models to directly generate document identifiers (DocIDs) (Li et al., 2024), which has aroused much attention. By memorizing the documents as the parametric knowledge of LLM, this kind of method breaks the limitations of traditional IR in terms of document granularity and simple relevance matching (Nguyen & Yates, 2023), offering more flexibility and creativity, thus better meeting practical needs. However, two severe problems hinder the current LLM-based retrieval. 1)Feature Locality: LLM-based retrieval normally adopt language models to learn the mapping from queries to the relevant document DocIDs. However, these DocIDs actually can not fully represent the global information of the passages. Meanwhile, directly feeding the whole passages into LLM is costly and infeasible, causing a trade-off between effectiveness and efficiency. 2)Retrieval Variance: Due to the inherent generative inconsistency property of large language models (Fan et al., 2024), current LLM-based retrieval may generate unforeseeable variances, especially when the input query or the size of context varies, directly results in undesirable and vulnerable performance which may not be preferred.
@@
Considering the problems mentioned above and better leveraging the capability of LLMs, we propose a fully LLM-based architecture with a two-stage fine-tuning method called Invar-RAG, as illustrated in Figure 1. In the retrieval stage, our approach initializes the pre-trained LLaMA (Yao et al., 2014) as the backbone and follows the bi-encoder architecture in DPR (Karpukhin et al., 2020) to construct our retriever. Compared to normal GR methods which need an iterative process of encoding and decoding, we introduce a component called LLM-aligned Retrieval. It first represents the input query and corpora into high-dimension space using a small language model (MiniLM) (Wang et al., 2020), then introduces a new loss function constructed by KL-divergence to align the coarse query-documents pairs representation to the LLM’s representation space. This allows the retriever to leverage the rich prior knowledge of LLM, typically addressing the feature locality caused by only feeding DocIDs to LLM. Moreover, based on the initial objective of our LLM-aligned Retrieval, we introduce the invariance loss to overcome the variance in the retrieval stage. By recognizing the invariant pattern that contributes the most to the performance and gradually forcing the model to rely on the invariant pattern, we can avoid the unforeseeable variances in practice and enhance the robustness of our RAG system. Finally, in the generation stage, we freeze the weights we fine-tuned before and optimize the generation function to allow the LLM to give correct answers to the retrieved documents. Our contributions are summarized as follows:
@@
Figure 1. Overview of proposed Invar-RAG.
2. Methodology
@@
2.1. Overall Framework of Invar-RAG
@@
2.2. Retrieval Stage
Architecture. For the Retriever architecture, we follow the approach of previous work (Ma et al., 2024), using the bi-encoder architecture from DPR (Karpukhin et al., 2020), but replacing the backbone model with LLaMA (Yao et al., 2014). Considering the efficiency, we first compute the vector embedding of a document ${d}^{R}_{i}\in\mathcal{D}^{R}$ as:
(1) ${V}_{r}({{d}^{R}_{i}})=\operatorname{Decoder}\left(`t_{1}␣t_{2}␣\cdots␣t_{k}^{\prime}\right)\left[-1\right]$
@@ 2.2. Retrieval Stage
(2) $s_{raw}\left(V_{q},V_{r}({{d}_{i}^{R}})\right)$ $=V_{q}\cdot V_{r}({{d}_{i}^{R}})$
@@
LLM-aligned Retrieval. Building on the above structure, we enhance our LLM-based retriever’s ability to return more relevant documents. We introduce LLM-aligned retrieval with invariance loss in the retrieval stage, effectively addressing the aforementioned issues. Current alignment methods, such as RA-DIT (Lin et al., 2023b), primarily focus on aligning the scoring functions between the retriever and the generator. However, the initial structural differences between novel retrievers (e.g., DRAGON+ (Lin et al., 2023a)) and LLMs still impede further optimization of the overall RAG system. Therefore, we design a novel LLM-based retriever to resolve this issue. Unlike previous LLM-based GR methods (Li et al., 2024), we no longer need to use DocIDs to retrieve relevant documents, which may cause the feature locality issue mentioned earlier. Instead, we adopt a fine-tuned alignment process that enables the LLM to perform representation learning. We leverage LoRA architecture (Hu et al., 2021) to add additional adapter parameter $\theta_{\mathcal{R}}$ to our raw representation $V_{r}$, denoted as $V_{r}({{d}_{i}^{R}},\theta_{\mathcal{R}})$. The corresponding relevance score can then be re-normalized among top-k relevant chunks $\mathcal{D}^{{R}^{\prime}}\subset\mathcal{D}^{{R}}$ as:
(3) $S^{r}_{\mathcal{R}}(V_{r}({{d}_{i}^{R}},\theta_{\mathcal{R}})|V_{q})=\frac{\exp{s_{raw}\left(V_{q},V_{r}({{d}_{i}^{R}},\theta_{\mathcal{R}})\right)}}{\sum_{{d_{i}^{{R}}}^{\prime}\in\mathcal{D}^{{R}^{\prime}}}\exp{s_{raw}\left(V_{q},V_{r}({{{d}_{i}^{R}}^{\prime}},\theta_{\mathcal{R}})\right)}}$
@@
For each document in the corpus, we need to compute $S^{r}_{\mathcal{R}}$ for $n$ times ($n$ represents the number of documents in $\mathcal{D}^{R}$) to rank the relevance scores. Consequently, the initial loss function for representation learning can then be defined by minimizing the KL-divergence (Kim et al., 2021) of two relevance scores leveraging Eq. 3.
(4) $\mathcal{L}_{rl}\left(\mathcal{D}^{R}\right)=\mathbb{E}_{d_{i}^{R}\in\mathcal{D}^{R}}\operatorname{KL}(S^{r}_{\mathcal{R}}(V_{r}({{d}_{i}^{R}},\theta_{\mathcal{R}})|V_{q})\|S^{r}_{\mathcal{R}}(V_{p}({{d}_{i}^{R}})|V_{q}))$
@@
Following the previous works (Lin et al., 2023b; Ma et al., 2020), fine-tuning both encoders hurt the performance (Bao et al., 2024), we only update a part of our initialized retriever, which is in charge of computing the query representation.
@@
Invariance Loss. To further enhance retrieval accuracy while maintaining robustness, we introduce invariance loss, building on our initial KL-divergence loss. Current refinement methods primarily rely on query rewriting (He et al., 2016; Chan et al., 2024) or LLM generation (Fan et al., 2024) to expand the search space and re-rank document chunks. However, they fail to recognize the effectiveness of different rewriting and generation procedures, directly resulting in the invariance problem. Specifically, we begin by rewriting the query and adjusting the context window to broaden the search space. Since not all refinement methods are effective, we identify invariant patterns to preserve retrieval performance while gradually incorporating weighted variant patterns to broaden the search space. To achieve this, we use the LSR score from LM-Supervised Retrieval (Shi et al., 2023) to determine whether a document effectively enhances the LLM’s answer prediction capability. For a training sample $(q,y)$, where q and y respectively represent the input query and output result, we first define the output probability of LM as:
(5) $p_{LM}(y|{V}(d_{i}^{R}\circ x))=\sum_{{d}^{R}_{i}\in\mathcal{D}^{R}}p_{LM}(y|{V}({d}^{R}_{i}\circ x))\cdot P_{\mathcal{R}}({d}^{R}_{i}|x)$
@@
(6) $P_{LSR}(d_{i}|q,y)=\frac{\exp(p_{LM}(y|d_{i}\circ q)/\tau)}{\sum_{d_{i}^{\prime}\in{\mathcal{D}^{R}}^{\prime}}\exp(p_{LM}(y|d_{i}^{\prime}\circ q)/\tau)}\approx\frac{\exp(p_{LM}(y|d_{i}\circ q)/\tau)}{\sum_{d_{i}^{\prime}\in\mathcal{D}^{R}}\exp(p_{LM}(y|d_{i}^{\prime}\circ q)/\tau)}$
@@
(7) $\mathcal{L}_{invar}(\mathcal{D}_{in})=\operatorname{Var}_{\mathcal{D}\subseteq\mathcal{D}_{var}}(\mathbb{E}_{d_{in}^{R}\in(\mathcal{D}_{in}\cup\mathcal{D})}\operatorname{KL}(P^{r}_{\mathcal{R}}(V_{r}(d_{in}^{R},\theta_{\mathcal{R}})|V_{q})\|P^{p}_{\mathcal{R}}(V_{p}(d_{in}^{R})|V_{q})))$
@@
(8) $\min_{\theta_{R}}\mathcal{L}_{rl}+\lambda\mathcal{L}_{invar}$
@@
2.3. Generation Stage
In the generation stage, We followed the same architecture as in the retrieval stage for answer prediction. To improve the generative capability of LLM for leveraging the retrieved information better, followed by prior works (Lin et al., 2023b; Shi et al., 2023), we adopt another LoRA adapter to fine-tune our model on different tasks. Specifically, for the same training sample $(x,y)$, we retrieve the top-$\overline{k}$ relevant document chunks $\mathcal{D}_{G}^{\prime}\subset\mathcal{D}_{G}$ by performing our model on retrieval task. For each retrieved chunk $d_{i}\in\mathcal{D}_{G}^{\prime}$, we design a special fine-tuning example by prepending it to the prompt as background information and create k independent instances for one original example: $\left\{(d_{i}\circ x,y)|i=1,\cdots,\overline{k}\right\}$. Then, following the previous work (Lin et al., 2023b; Qi et al., 2020), we fine-tune the language model using the next-token prediction objective and minimize the loss as follows:
(9) $\mathcal{L}(\mathcal{D}_{G}^{\prime})=-\sum_{i}\log p_{LM}(y|d_{i}\circ x)$